A telephone company wants to create as
many 7-digit phone numbers as possible
without changing the first 3 digits. How
many phone numbers can be created?

Found 2 solutions by nyc_function, sudhanshu_kmr:
Answer by nyc_function(2741)

7P3 = 7!/(7 - 3)!

7P3 = 5040/4!

7P3 = 5040/24

7P3 = 210 phone numbers can be created.

Answer by sudhanshu_kmr(1152)


According to question first 3 digits are fixed.
Number of ways to create 4 digit numbers = 10* 10*10*10 = 10000
(repetition of digits are allowed on phone number)

if any doubt, feel free to contact....